A) Suppose X is normally distributed with mean 10 and standard deviation 2. Then P(X > 15) = P(Z > __) where Z is standard normal. (Fill in the blank.)
Answer #1

Solution :

Given that,

mean = = 10

standard deviation = = 2

P(x > 15) = 1 - P(x< 15)

= 1 - P[(x -) / < (15-10) /2 ]

= 1 - P(z <2.5 )

Using z table

= 1 -  0.9938

probability=0.0062

answer (Fill in the blank.)

P(Z > _2.5_) where Z is standard normal.